Banko Inc. manufactures sporting goods. The following information applies to a machine purchased on January 1, Year 1: Purchase price $ 91,000 Delivery cost $ 5,000 Installation charge $ 3,000 Estimated life 5 years Estimated units 160,000 Salvage estimate $ 3,000 During Year 1, the machine produced 56,000 units and during Year 2, it produced 58,000 units. Required Determine the amount of depreciation expense for Year 1 and Year 2 using each of the following methods:

Straight-line

double-declining-balance

units of production
Answer #1

Total Cost of Machine = Purchase price + Delivery cost + Installation charge

= 91000+5000+3000 = $99,000

Depreciation expense for Year 1 (Straight-line) = (Cost - salvage value) / estimated life of the assets

= (99000-3000)/5 = $19,200Depreciation expense for Year 2 (Straight-line) = $19,200

Double declining depreciation rate = Straight line depreciation rate x 2

= (100/5 years) x 2

= 40%

Depreciation expense for Year 1 (double-declining-balance) = 99000*40% = $39600

Depreciation expense for Year 2 (double-declining-balance) = (99000-39600)*40% = $23,760

Depreciation expense per unit (units of production) = (99000-3000)/160000 = $0.60 per unit

Depreciation expense for Year 1 (units of production) = 56000*.60 = $33600

Depreciation expense for Year 2 (units of production) = 58000*.60 = $34800